# Mixtera: A Data Plane for Foundation Model Training

### v2

@@ 4. Mixtera’s Design
We propose to address these challenges for training data management by building Mixtera, a foundation model training data plane. We derive the following design goals for such a system.
@@
Goal G2: The system needs to be lightweight, i.e., not require many components to set up and be easily integratable into existing training setups.
@@
Goal G3: While being user-friendly and flexible, the system needs to ensure high-throughput and determinism (reproducibility).
@@
Goal G4: The system must support adjusting the mixture dynamically during the training.
4.1. System Overview and Design
TODO!
@@ 4.1. System Overview and Design
Inspired by the Lakehouse architecture (Armbrust et al., 2021), we design Mixtera as a read-only layer that can be deployed on top of existing training data collections, which are typically stored in a distributed filesystem. Figure 3 shows the architecture of the system. Mixtera manages a centralized database of metadata (i.e., properties such as language or source dataset) of all training data samples (G1). A sample can be a piece of text (for LLM training) or a text-image pair (for VLM training). Mixtera assigns every sample a unique ID and properties instead of treating the training data as a blob of homogeneous, contiguous data. It allows users to declaratively query the relevant samples for a training. To be lightweight (G2), Mixtera does not reorganize or modify the data files on disk. For model training, it provides a standard iterator that can be used, e.g., in conjunction with a torch.DataLoader. Mixtera is agnostic to the model training framework, supports training interruptions using checkpoints, and ensures determinism (G3) through careful shuffling, i.e., for identical queries, Mixtera always provides data in an identical order, which is important for reproducibility as well as for debugging issues like loss spikes (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025). It supports adjusting the mixture during training (G4) by transferring chunks (lists of pointers to samples) whose mixture can change over time.
[ frame=lines, framesep=2mm, baselinestretch=0.8, bgcolor=LightGray, fontsize=, linenos ] python client = MixteraClient("127.0.0.1", 8080) job_id = "test_job" query = Query.for_job(job_id).select(("license","==","CC")) mixture = StaticMixture( MixtureKey("language": ["JavaScript"]): 0.7, MixtureKey("language": ["HTML"]): 0.3 , chunk_size=1024) qea = QueryExecutionArgs(mixture=mixture, num_workers=4, dp_groups=1, nodes_per_group=1) rsa = ResultStreamingArgs(node_id=0, dp_group_id=0, job_id=job_id) ds = MixteraTorchDataset(client, query, rea, rsa) dl = torch.utils.data.DataLoader(ds, batch_size=1024, num_workers=4)
[…]
Query interface. In Figure 4, we show an example of a query that statically selects only Creative Commons data, and then mixes HTML and JavaScript data in a 70:30 ratio during training. Mixtera’s implementation takes care of executing the query and obtaining the samples without needing to worry about correctness, even in distributed training. The user only needs to provide the ID of the node and its data parallel group, which is obtained from the training framework. Mixtera currently allows to express static filter operations on properties, and static as well as dynamic mixtures across all properties (G4). For more details on supported mixture types, we refer to Section 5.2.2.
Dataflow. Mixtera follows a server-client model. As shown in Figure 3, the server runs on a node and each training node runs client instances. Users first register samples at the server to populate the metadata database. They then can submit queries. A query is executed at the server in two phases. First, Mixtera applies static filters from the query (e.g., English-only) to obtain all samples eligible for training. Second, during training, the server distributes chunks of that query result to the client(s), which specify which samples to train on. The server ensures that the chunks are distributed correctly, i.e., tensor- and pipeline parallel stages receive the same input data. The server generates chunks according to the current mixture, i.e., it iteratively takes samples from the query result generated in the first phase of query execution, such that data in the chunk follow the current mixture. As an iterable data loader, Mixtera faces the challenges of determinism and checkpointing. We address this by shuffling based on the query and support to load/store the query state.
Chunks. Mixtera does not store the sample payloads, but rather only the metadata of each sample. During training, the nodes receive chunks. A chunk is a fixed-size collection of pointers to samples in files. They tell the client which samples which file to load and train on (e.g., sample 10 in file wikipedia.jsonl.zst). The files can be local, in a distributed filesystem, or cloud storage.
@@
Storing and distributing pointers to samples instead has several advantages. First, users can store data in their locations of choice (e.g., an object store, or an distributed filesystem). Chunks are independent of the filesystem structure (G1). Second, it allows Mixtera to support dynamic mixtures (G4), as the data composition of chunks can change over time. Third, the pointer-model avoids creating a data fetching bottleneck at the Mixtera server. The server only creates chunks and each client fetches the data they need. Fourth, we avoid a lock-in effect on Mixtera and allow for easy adoption on existing data collections (G2). Last, we natively support other modalities like images, which would not be straightforward to ingest at scale into a database.
Mixtures. Users can use any subset of sample properties to define a mixture. Users can even change the properties that they mix on during training with dynamic mixture schedules. To handle this, Mixtera implements a MixtureKey abstraction that allows users to describe which samples they want to use. The keys also implement flexible property matching to define which samples to use during chunk generation (Section 5.2.2).
@@
High-throughput data fetching. The challenge of using chunks is that we need to handle suboptimal data layout. Files may have arbitrary property distributions, or might follow a partial structure (e.g., a file only contains Wikipedia data, but the languages are distributed randomly). We cannot re-organize data within the existing files. Formats like jsonl were not built with random access in mind, but chunks force clients to load individual samples from files. To avoid data stalls (G3), we implement chunk generation to use subsequent sample ranges in files if the samples have the same properties and engineer Mixtera’s implementation to fetch those ranges as efficiently as possible (Section 5.3).
@@
Open source. Mixtera comes as a Python package that provides the entrypoint for the server and abstractions for the client. Its codebase, consisting of approximately 11 k lines of Python and C++ (excluding tests), is open-source111Available at https://github.com/eth-easl/mixtera.. It is rigorously tested with a full set of unit and integration tests. We are continuously working on enhancing the system and welcome contributions.
@@ 5. Implementation
We explain how sample metadata is managed (Section 5.1) at the server, how it executes queries and creates chunks (Section 5.2), how those chunks are parsed at the client (Section 5.3), and give details on Mixtera’s integration into training frameworks (Section 5.4).
5.1. Sample Management
Mixtera manages samples in a MixteraDataCollection (MDC). It keeps track of all samples and their properties. The MDC leverages DuckDB (Raasveldt and Mühleisen, 2019) as the robust and flexible metadata management system underlying it. Every registered sample has properties according to a schema defined by a MetadataParser, e.g., language. Representing the structure of training data, Mixtera initializes the database with three tables for source datasets, files, and samples. The source dataset can also be ignored and stored as a schema property instead, to support collections such as The Pile (Gao et al., 2020) where data from different datasets is contained within the same files.
MetadataParsers. A MetadataParser is a class that defines a schema for a data collection and, given a sample, extracts and returns the properties according to the schema. A schema is a list of properties which have a type (e.g., string or enum), a nullable field, and a multiple field, describing whether a single sample can take multiple values for this properties, e.g., multiple languages. Mixtera maps this Python schema to a proper database schema, comes with a set of pre-defined parsers for common datasets, and enables users to define custom parsers. For example, there is a parser for The Pile (Gao et al., 2020). Every sample has the property pile_set_name describing the source dataset. This is a non-nullable, non-multiple, enum property, since we know the limited set of possible values.
Data types. Based on the parser, Mixtera adjusts the DuckDB sample table. Properties marked as multiple get mapped to DuckDB’s list type, e.g., a multiple string property gets mapped to VARCHAR[]. Enum properties are appropriately inserted into DuckDB. Using enum instead of strings for properties where all values are known beforehand speeds up query execution in DuckDB.
Insertion. When inserting (registering) data, the MDC accumulates all relevant files, and prepares the DuckDB table schema. Then, all files are inserted into the files table, and a pool of workers in parallel processes all files using the MetadataParser to extract the sample metadata. We then aggregrate the worker results and insert it into the database, as DuckDB does not support insertion from multiple Python process workers. We found that converting the worker results to columnar pyarrow in-memory tables and then inserting into the samples table has the highest throughput.
@@ 5.2. Server-Side Query Execution
After receiving a query from the client, the server executes it in two phases. The first phase is performed via the MDC DuckDB-abstraction and applies static filters to identify all potentially relevant samples, e.g., all non-English samples are filtered out, and groups consecutive samples into intervals. The second phase constructs a specialized data structure called ChunkerIndex that enables efficient, mixture-aware chunk generation.
@@ 5.2.1. SQL generation and interval detection
After receiving an object representation of the query (c.f. Figure 4), similar to ORM frameworks like sqlalchemy, the Mixtera server generates a base SQL query from this object. This query returns a table in which each row represents a sample that the user is interested in. Mixtera ensures that the generated SQL matches the MDC’s table schema, e.g., whether a property can have multiple values or not.
@@
Even though samples 4-5 (file 1) and 1 (file 2) share the same properties (Python, Apache), they are in different files and thus form separate intervals. The primary key is formed by the sample and file ID together. Mixtera constructs a SQL query that processes the data in multiple stages:
First, Mixtera establishes a Common Table Expression (CTE) named base_data that contains the filtered samples: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql WITH base_data AS ( – Our generated base filtering query here, e.g., SELECT * FROM samples WHERE license = ’MIT’),
Next, Mixtera identifies breaks in the sample sequence using window functions. The grouped_samples CTE calculates the difference between consecutive sample IDs within groups sharing the same properties: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql grouped_samples AS ( SELECT *, sample_id - LAG(sample_id, 1, sample_id) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS diff FROM base_data), Here, a diff value of 1 indicates consecutive samples, while any other value indicates a break in the sequence.
The intervals CTE then groups the sequences into intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql intervals AS ( SELECT file_id, lang, license, SUM(CASE WHEN diff != 1 THEN 1 ELSE 0 END) OVER (PARTITION BY file_id, lang, license ORDER BY sample_id) AS group_id, MIN(sample_id) as int_strt, MAX(sample_id)+1 as int_end FROM grouped_samples GROUP BY file_id, lang, license, diff, sample_id) The group_id is incremented when there is a break in the sequence, creating unique identifiers for each interval.
Finally, Mixtera aggregates the results to get the final intervals: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]sql SELECT file_id, lang, license, group_id, MIN(int_strt) as interval_start, MAX(int_end) as interval_end FROM intervals GROUP BY file_id, lang, license, group_id ORDER BY file_id, interval_start;
While we initially implemented the interval aggregation within Python, letting DuckDB optimize and parallelize the calculation of the interval is more efficient. This optimization can only improve I/O if samples within files are clustered by properties and not randomly distributed. For example, if a user has a file structure based on the source dataset and mixes based on that, Mixtera can heavily leverage the range optimization and just read, e.g., all Wikipedia samples from the sample file. If the user decides to now mix on language as well, there might only be local clusters within the files. Since Mixtera is read-only by design, it cannot re-shuffle data.
@@ 5.2.2. Chunk generation
The chunk generation algorithm is based on the ChunkerIndex data structure, which organizes sample ranges by their properties. This data structure is generated based on the output from the interval query. Before explaining its construction, we first introduce the key abstraction that enables flexible property matching: MixtureKey.
MixtureKey abstraction. A MixtureKey represents a set of properties and their values, e.g., language being English and license being MIT. A property in a key can have multiple values. A key matches another if their lists of values intersect for all properties they share. This matching is crucial because the resulting interval table from DuckDB contains the full cross-product of all properties—a sample might have values for language, license, size, topic, and more—while a mixture specification typically considers only a subset of these properties. For example, a mixture might specify language:English to select English text regardless of license, while the underlying samples are stored with full property information (e.g., language:English,German;license:CC and language:English;license:MIT). The flexible matching enables finding all samples that satisfy certain properties while ignoring irrelevant attributes they might have. It also allows us to define mixtures on multiple properties with multiple values, instead of being limited to a single property (c.f. Section 3). To ensure deterministic behavior, we implement a total ordering over keys based on the number of properties, property names, and their values. We sometimes refer to a specific MixtureKey as a domain, e.g., the key for language:English defines the domain for English samples.
The ChunkerIndex. The ChunkerIndex is a nested mapping from MixtureKeys to sample locations. We refer to keys in the ChunkerIndex as component keys. For each unique combination of all available properties and values, the index maps to dataset IDs, which in turn map to file IDs and finally to sorted lists of intervals of samples which share the properties and values as specified in the MixtureKey. While the index maintains the complete property information of samples, the matching semantics allows us to efficiently query arbitrary subsets of these properties. Consider a simplified example where we represent MixtureKeys as strings. A fragment of the ChunkerIndex might look like: {minted}[frame=lines,framesep=2mm,baselinestretch=0.8,bgcolor=LightGray,fontsize=]python "language:JavaScript,HTML;license:MIT": ds_1: file_1: [(1,4), (10,15)], # half (right) open ranges file_2: [(1,2)] , "language:Python;license:Apache": ds_1: file_1: [(1,2)]
@@
Building the ChunkerIndex. The index is built in parallel in a C++ extension, processing the interval table from DuckDB provided in Apache Arrow format. Operating on the Arrow table in Python would be too slow due to Global Interpreter Lock (GIL) constraints. Using multiprocessing to circumvent the GIL would require expensive pickling of nested dictionaries. Our C++ implementation uses multithreading and only acquires the GIL at the end when creating the final Python result object.
@@
Chunk generation. A mixture in Mixtera is represented as a mapping from MixtureKeys to their target proportions within a chunk. Given a chunk size that describes how many samples each chunk should contain, these proportions are converted to absolute counts following the largest remainders quota method. Users can set a mixture to be strict, requiring exact proportions, or best-effort, allowing to continue to generate chunks even if the mixture cannot be exactly fulfilled. The chunk generation algorithm (Algorithm 1) iteratively constructs chunks that conform to the current mixture. This algorithm supports dynamic mixture, as the mixture can be changed between chunks.
Similar to the ChunkerIndex, a chunk is also structured as a hierarchical dict with keys and files pointing to lists of ranges. This chunk can be parsed on a training node by loading the sample ranges from the files (Section 5.3). While the ChunkerIndex uses all properties/values as MixtureKeys (component keys), a chunk’s keys are identical to the mixture’s keys, which might only contain a subset of properties.
@@
For each key in the mixture, the algorithm keeps track of how many samples we still need to put into the chunk that is currently being generated (remaining_sizes). For each key in the mixture (line 6), it checks whether it matches a component key in the chunker index (lines 7-8). If we find a match, we try and obtain samples (ranges) from the ChunkerIndex for this component key (lines 9+).
A call to obtain samples for a component key can return fewer samples than requested, e.g., if we are looking for JavaScript data and we need 5 samples, but we only have 3 JavaScript/MIT licensed samples, the according component key can only return 3 samples. Requesting samples is implemented as requesting ranges (from potentially multiple files) such that the overall number of samples in the returned list of ranges is . The lists of ranges per file are merged into the existing sorted list of ranges. Mixtera’s implementation uses Python generators that yield ranges of samples and accept the number of samples needed as input through the Python generator’s send mechanism. This hides the complexity of the take samples operation. It also allows for efficient, stateful iteration over the available ranges while maintaining control over how many samples are returned in each request. There is one generator per component key that returns the ranges containing samples based on the ChunkerIndex but ensures ranges are split such that never too much data is returned.
@@
If after traversing all component keys we did not find sufficient samples for a key in the mixture, in strict mode, chunk generation fails. In best-effort mode, the algorithm redistributes any unfulfilled counts to the remaining mixture components proportionally to their original ratios. For example, if we need 100 JavaScript samples but only find 80, the remaining 20 samples would be proportionally distributed among other components. To avoid infinite loops, we only distribute samples to keys on which we were able to find any samples in the last iteration.
@@
Mixtera’s implementation of this algorithm ensures determinism because (1) the mixture’s keys are processed in a consistent order and (2) when multiple component keys match a mixture key, they are considered in a deterministic order based on a seeded shuffle of all possible keys. This approach ensures that identical queries with identical mixtures always produce identical chunks, important for training debugging and reproducibility (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
Mixtera also supports an alternative, arbitrary chunk generation algorithm. It does not require any specified mixture and builds up chunks by iterating over the component keys, moving on to the next one when one key is depleted. This can be useful if the data is completely property-free and guarantees maximum throughput as all ranges are consecutive.
@@
Static Mixture: Users explicitly specify fixed proportions for different property combinations (Figure 4). This supports arbitrary properties and is not limited by, e.g., directory boundaries.
@@
Hierarchical Mixture: An advanced static mixture that allows to specify nested property relationships. For example, users can define that 50 % of the data should be legal texts, and within that, 60 % should be in English and 40% in French. Mixtera automatically flattens this hierarchy into appropriate MixtureKeys.

### v3

@@ 4. Mixtera’s Design
We address these challenges for training data management by building Mixtera, a foundation model training data plane. We derive the following design goals for such a system.
@@
Goal G2: The system needs to be lightweight, i.e., easily integrate into existing training deployments without requiring setting up many components.
@@
Goal G3: The system needs to ensure high-throughput, determinism, and reproducibility, while being user-friendly and flexible.
@@
Goal G4: The system must support adjusting the mixture dynamically during training.
4.1. Data Model
Mixtera operates on training datasets of the following structure:
Sample. A sample is the atomic unit of training data, as supplied by the user. In LLM training, the sample unit is typically a text string or a pre-tokenized sequence; VLM training uses multimodal samples (e.g., text-image pairs). Sample granularity varies in practice, ranging from short texts to full documents.
Property. A property is a named attribute of a sample, with a domain of possible values. Properties characterize samples and can be single-valued (each sample has exactly one value, e.g., ), or multi-valued (each sample may have multiple values, e.g., ).
Data Collection. A data collection is a tuple where is a finite set of samples, is a set of properties, and maps each (sample, property) pair to a subset of values from that property’s domain. We also refer to it as the sample metadata.
Mixtera defines mixtures fully independent of the filesystem–it is fully based on (logical) properties.
Mixture Key. A mixture key is a partial mapping from a subset of properties to sets of values, i.e., , with . For mixture keys and , we say matches , if f.a. , . This relation is non-symmetric.
Component Key. A sample induces a component key , s.t. .
Data Mixture. A data mixture is a mapping from mixture keys to target proportions: with .
4.2. System Overview
Mixtera is a read-only layer that can be deployed on top of existing sets of samples , which are typically stored in a distributed filesystem or cloud object store. Figure˜3 shows the client-server architecture of the system. The server runs on one node, and each training node runs client instances. Mixtera manages a centralized database which stores the mapping , i.e., the sample metadata (G1). This database needs to be populated by scanning all samples once before it can be used for training. Since the metadata typically consists of integers and floats, it is much smaller than the actual samples. Mixtera defers the reading of actual sample payloads to clients. This enables to scale to large-scale training datasets.
@@ 4.2. System Overview
Mixtera assigns every sample a unique ID. It allows model developers to declaratively query the relevant samples for a training job. To remain lightweight (G2), Mixtera does not reorganize or modify the data files on disk. It provides a standard iterator for model training that can be used in conjunction with loaders like torch.DataLoader. Mixtera is agnostic to the training framework, supports training interruptions using checkpoints, and ensures determinism (G3) through careful shuffling, i.e., for identical queries, Mixtera always provides data in an identical order, which is important for reproducibility and debugging issues like loss spikes (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025). It supports adjusting the mixture during training (G4) by transferring chunks (lists of pointers to samples) whose mixture can change over time.
Data types. Mixtera supports diverse sample types (e.g., pre-tokenized text in binary files, strings in jsonl/ parquet, text–image pairs in webdatasets). This covers many relevant use cases. Even modern foundation models for tabular data treat tables as strings (Bodensohn et al., 2025). The sample granularity is inherent to the ingested dataset. As Mixtera is a read‑only layer, it adopts whatever granularity the data already uses. While some features such as tokenization or image decoding are modality-specific, the system is fundamentally agnostic to the sample unit.
Chunks. Chunks are Mixtera’s core abstraction for scheduling and mixture enforcement. Abstractly, a chunk is a collection of samples of fixed size , and the set conforms to the current mixture . That is, for each mixture key , the proportion of samples in matching is approximately . In implementation, Mixtera does not materialize or transfer the sample payloads . Instead, chunks contain only pointers to samples: metadata specifying which samples in which files to load (e.g., “samples 100-150 in wikipedia.jsonl.zst”).
@@
Deferred reading. Storing only metadata and distributing pointers to samples (as opposed to ingesting the actual sample payloads into the system) has several advantages. First, users can store data in their locations of choice (e.g., an object store, or a distributed filesystem). Chunks are independent of the filesystem structure (G1). Second, it allows Mixtera to support dynamic mixtures (G4), as the data composition of chunks can change over time. Third, the pointer-model avoids creating a data fetching bottleneck at the Mixtera server. The server only creates chunks and each client fetches the data they need. Fourth, we avoid a lock-in effect and allow for easy adoption on existing data collections (G2). Last, we natively support other modalities (whereas image or video payloads would not be straightforward to ingest at scale into a database).
Query interface. In Figure˜4, we show an example query that statically selects only Creative Commons data, and then mixes HTML and JavaScript data in a 70:30 ratio during training. Mixtera takes care of executing the query and obtaining the samples without needing to worry about correctness, even in distributed training. The user only needs to provide the ID of the node and its data parallel group, which is obtained from the training framework.
Mixtera allows expressing static filter operations on properties via SPJ-style predicates, and static as well as dynamic mixtures across all properties (G4). The static filter defines the ground set of all potential data (e.g., only CC), while the mixture describes the proportions of how the data from this set is mixed during training (JavaScript:HTML ratio). The MixtureKey class operationalizes its formal definition (Section˜4.1) in a user-friendly Python API.
Executing a query. Before submitting a query, users catalog all samples at the server. This involves defining the data schema as a Python class, which allows Mixtera workers to scan all samples and populate the database. Users then can submit queries. A query is executed at the server in two phases. First, Mixtera applies static filters from the query (e.g., English-only) to obtain all samples eligible for training (QueryResult). Second, during training, the server distributes chunks of that query result to the client(s), which specify which samples to train on. The server ensures that the chunks are distributed correctly, i.e., tensor- and pipeline parallel stages receive the same input data. The server generates chunks according to the current mixture, i.e., it iteratively takes samples from the query result such the chunk abides by the current mixture. As an iterable data loader, Mixtera faces the challenges of determinism and checkpointing. We address this by shuffling based on the query and support to load/store the query state.
@@
High-throughput data fetching. The challenge with not re-organizing the user’s data files is that Mixtera needs to handle suboptimal data layout. Files may have arbitrary distributions of data properties (e.g., even if a file only contains Wikipedia data, data in different languages might be distributed randomly across files). Formats like jsonl were not built with random access in mind, yet chunks force clients to load individual samples from files. To avoid data stalls (G3), Mixtera reads subsequent samples in files if the samples follow the same properties (e.g., same language). This is achieved through intervals identifying consecutive samples in file that share the same component key . A chunk is thus represented as a collection of intervals whose total size equals and whose union satisfies the mixture. We use nested CTEs in DuckDB to find these intervals as fast as possible (Section˜5.3). Users are not required to use sequential file formats.
Alternative design considerations. A metadata SQLite database at each client might seem like a simple alternative. While DBs can federate and query external sources with SPJ-style predicates, we still need an interface between the storage, query engine, and training framework, which can turn query results into input tensors (G1, G2). Furthermore, using a database at each client would not enable dynamic mixing algorithms (G4). Mixtera serves as the interface that connects a metadata database to the training framework, with a full feature set for dynamic model training. It also alleviates ML engineers from manually maintaining the database. We do not find centralized metadata management to be a bottleneck even for large-scale training jobs, because the work at the service is minimal (occasional chunk creation and distribution). Mixtera maintains high training throughput for large training jobs (Section˜6.3). Industry frameworks like Google’s Pathways (Barham et al., 2022) also rely on a single Python process to coordinate large-scale training. A current limitation of Mixtera’s design is that it does not enable incorporating new data into an existing query. However, this (i) affects nearly all dataloaders as they all build some internal state, and (ii) and can in all cases be dealt with by executing a new query, e.g., in continual pretraining from a model checkpoint.
@@
Open source ecosystem. Mixtera comes as a Python package that provides the entry point for the server and abstractions for the client. The codebase, consisting of approximately 11 k lines of Python and C++ (excluding tests), is open-source111Available at https://github.com/eth-easl/mixtera.. It is rigorously tested with a full set of unit and integration tests. We are continuing to add features and welcome contributions.
@@ 5. Implementation
We explain how Mixtera ingests sample metadata (Section˜5.1), how it executes queries and creates chunks (Section˜5.2), how those chunks are parsed at the client (Section˜5.3), and describe Mixtera’s integration into training frameworks (Section˜5.4).
5.1. Metadata Insertion
Mixtera manages the metadata in an MixteraDataCollection (MDC) which uses DuckDB (Raasveldt and Mühleisen, 2019) as the underlying DBMS to efficiently store and query sample properties. The implementation itself is agnostic to the DBMS. While DuckDB’s pluggable extension architecture (or similar approaches like PostgreSQL’s Foreign Data Wrappers or SQLite’s Virtual Tables) would allow to query the sample files directly, to defer reading (Section˜4.2) to the training nodes, the MDC serves as a metadata index.
Initial ingestion. To populate the MDC, users need to define a MetadataParser. A MetadataParser operationalizes the property set and extraction of for each sample . It defines the schema specifying and domains for each property. In Python, the schema is a list of properties which have a type (e.g., string or enum), a nullable field, and a multiple field, describing whether a single sample can take multiple values for this properties (e.g., several languages). Mixtera adjusts the underlying database table by mapping the Python schema to a proper database schema. The system comes with a set of pre-defined parsers for common datasets and enables users to define custom parsers.
File scanning. Metadata ingestion is a one-time preprocessing step. The MDC first accumulates all data files and prepares the database schema. Mixtera then parallelizes metadata extraction using a worker pool. Workers process files in batches, where each worker sequentially reads samples from its assigned files and applies the MetadataParser to extract property values (e.g., parsing JSON fields). Since DuckDB does not support concurrent insertions from multiple processes, Mixtera aggregates worker results in the main process. To optimize insertion throughput, we convert the collected metadata to columnar PyArrow in-memory tables before bulk insertion into DuckDB. Importantly, metadata extraction is decoupled from training. Users must define properties upfront, but can later add new properties by re-scanning the dataset with an updated parser without affecting existing metadata.
@@ 5.2. Server-Side Query Execution
After registering data, users can execute queries. When a client sends a query, the server executes it in two phases. The first phase is performed via the MDC and applies static filters to identify all relevant samples, and groups consecutive samples into intervals. The second phase constructs a data structure called ChunkerIndex that enables efficient, mixture-aware chunk generation.
@@ 5.2.1. SQL generation and interval detection
After receiving an object representation of the query (c.f. Figure˜4), similar to ORM frameworks like sqlalchemy, the Mixtera server generates a base SQL query from this object. This query returns a table in which each row represents a sample that the user is interested in. Mixtera ensures that the generated SQL matches the MDC’s table schema, e.g., whether a property can have multiple values or not.
@@
Even though samples 4-5 (file 1) and 1 (file 2) share the same properties (Python, Apache), they are in different files and thus form separate intervals. The primary key is formed by the sample and file ID. Mixtera constructs a SQL query that processes the data in multiple stages:
First, Mixtera establishes a Common Table Expression (CTE) named base_data that contains the filtered samples:
WITH base_data AS (– Our generated base filtering query here, e.g.,SELECT * FROM samples WHERE license = ’MIT’), -
Next, Mixtera identifies breaks in the sample sequence using window functions. The grouped_samples CTE calculates the difference between consecutive sample IDs within groups sharing the same properties:
grouped_samples AS (SELECT *, sample_id - LAG(sample_id, 1, sample_id)OVER (PARTITION BY file_id, lang, licenseORDER BY sample_id) AS diffFROM base_data),Here, a diff value of 1 indicates consecutive samples, while any other value indicates a break in the sequence.
The intervals CTE then groups the sequences into intervals:
intervals AS (SELECT file_id, lang, license,SUM(CASE WHEN diff != 1 THEN 1 ELSE 0 END)OVER (PARTITION BY file_id, lang, licenseORDER BY sample_id) AS group_id,MIN(sample_id) as int_strt, MAX(sample_id)+1 as int_endFROM grouped_samplesGROUP BY file_id, lang, license, diff, sample_id)The group_id is incremented when there is a break in the sequence, creating unique identifiers for each interval.
Finally, Mixtera aggregates the results to get the final intervals:
SELECT file_id, lang, license, group_id,MIN(int_strt) as interval_start, MAX(int_end) as interval_endFROM intervalsGROUP BY file_id, lang, license, group_idORDER BY file_id, interval_start;
Using intervals of samples can only improve I/O if samples within files are clustered by properties and not randomly distributed. Since Mixtera is read-only by design, it does not re-shuffle data.
@@ 5.2.2. Chunk generation
After obtaining the query result with all relevant intervals, the server next runs the chunk generation algorithm. This algorithm is based on the ChunkerIndex data structure, which organizes sample ranges by their properties. We re-visit the MixtureKey concept from an implementation perspective.
MixtureKey abstraction. A MixtureKey represents a set of properties and their values. The class implements the formal definition from Section˜4.1: a partial mapping for some . The matching relation enables flexible querying: matches if . This matching is crucial as the resulting interval table from DuckDB contains the full cross-product of all properties—a sample might have values for language, license, size, topic, and more—while a mixture specification may consider only a subset of these properties, as we discuss in the next paragraph. It also allows us to define mixtures on multiple properties with multiple values, instead of being limited to a single property (c.f. Section˜3). To ensure deterministic behavior, we implement a total ordering over keys based on the number of properties, property names, and their values. We sometimes refer to a specific MixtureKey as a domain, e.g., the key for lang:English defines the domain of English samples.
The ChunkerIndex. Recall from Section˜4.1 that each sample induces a component key defined f.a. as . The ChunkerIndex organizes intervals by these component keys, i.e., , where is the filtered subset from the query, and intervals group consecutive samples sharing the same . While the index maintains the complete property information of samples, it enables efficient sample lookup: given a mixture key , we identify all component keys where matches , and retrieve their associated intervals. Consider a simplified example with MixtureKeys as strings. A fragment of the ChunkerIndex might look like:
@@
Building the ChunkerIndex. The index is built in parallel in a C++ extension, processing the interval table from DuckDB provided in Apache Arrow format. Operating on the Arrow table in Python would be too slow due to Global Interpreter Lock (GIL) constraints. Using multiprocessing to circumvent the GIL would require expensive pickling of nested dictionaries. Our C++ implementation uses multithreading and only acquires the GIL at the end.
@@
Chunk generation. As defined in Section˜4.2, a chunk is abstractly a collection of samples conforming to mixture . In implementation, contains only interval pointers rather than materializing samples. Given and chunk size , Algorithm˜1 constructs such that for each mixture key , the number of samples with component keys matching approximates . Users can set a mixture to be strict, requiring exact proportions, or best-effort (continue to generate chunks even if the mixture cannot be exactly fulfilled). The algorithm iterates through each , identifies all matching component keys in the ChunkerIndex, and extracts intervals until the target count is reached. This algorithm supports dynamic mixture, as the mixture can be changed between chunks.
@@
For each key in the mixture, the algorithm keeps track of how many samples we still need to put into the chunk that is currently being generated (remaining_sizes). For each key in the mixture (line 6), it checks whether it matches a component key in the chunker index (lines 7-8). If we find a match, we try and obtain samples (ranges) from the ChunkerIndex for this component key (lines 9+). A call to obtain samples for a component key can return fewer samples than requested, e.g., if we are looking for JavaScript data and we need 5 samples, but we only have 3 JavaScript/MIT licensed samples, the according component key can only return 3 samples. Requesting samples is implemented as requesting intervals (from potentially multiple files) such that the overall number of samples in the returned list of intervals is . The lists of intervals per file are merged into the existing sorted list of intervals.
@@
If, after traversing all component keys, we did not find sufficient samples for a key in the mixture, in strict mode, chunk generation fails. In best-effort mode, the algorithm redistributes any unfulfilled counts to the remaining mixture components proportionally to their original ratios. For example, if we need 100 JavaScript samples but only find 80, the remaining 20 samples would be proportionally distributed among other components. To avoid infinite loops, we only distribute samples to keys on which we were able to find any samples in the last iteration. This redistribution mechanism can be enabled or disabled, allowing users to either prioritize strict mixture fidelity (which may stop training when samples are exhausted) or training as long as possible with approximated proportions.
Implementation details. The take samples operation (line 9) is implemented using Python generators that yield ranges of samples and accept the number of samples needed as input through the generator’s send mechanism. This hides the complexity of range management and allows for efficient, stateful iteration over available ranges while maintaining control over sample counts. There is one generator per component key that returns ranges containing samples based on the ChunkerIndex, ensuring ranges are split such that excess data is never returned.
@@
Determinism. Mixtera’s implementation of this algorithm ensures determinism because (1) the keys are processed in a consistent order and (2) when multiple component keys match a mixture key, they are considered in a deterministic order based on a seeded shuffle of all possible keys. This ensures that identical queries with identical mixtures always produce identical chunks, which is important for debugging and reproducibility (Zhuang et al., 2022; Cooper et al., 2022; Qian et al., 2021; Karamcheti et al., 2021; Team, 2025).
Sampling and randomness. Chunks are generated randomly, but not completely i.i.d.: (i) each chunk must contain the proportions specified by the mixture , and (ii) all matching component keys are considered equal. Consider a mixture key that matches component keys and . The order in which and are processed is determined by a seeded shuffle, ensuring determinism across runs while providing randomization across different matching keys. Once a component key is selected, samples are drawn from its intervals. Within each file, intervals are used for I/O efficiency, but the order in which files are processed is shuffled.
Importantly, if the sample count for mixture key is satisfied by samples from , samples from may never be selected for that chunk. This is not a sampling bias but rather reflects the mixture specification. By omitting the license property from , the user declares all JavaScript samples equivalent regardless of license. To ensure unbiased sampling across all relevant dimensions, users should include all properties they care about in the mixture.
@@
Static Mixture: Users explicitly specify fixed proportions for different property combinations (Figure˜4). This supports arbitrary properties and is not limited by, e.g., directory boundaries.
@@
Hierarchical Mixture: An advanced static mixture that allows specifying nested property relationships. For example, users can define that 50 % of the data should be legal texts, and within that, 60 % should be in English and 40% in French. Mixtera automatically flattens this hierarchy into appropriate MixtureKeys.